OpenStudy (anonymous):

9. A farmer wished to construct a rectangular fence with one side along a barn. The farmer has 240 feet of fence and wishes to have the length be three times the width. If the side of the barn accounts for the length of the fence, find the dimensions of the fence.

OpenStudy (anonymous):

The dimensions of the fence are W x L L=3W It really depends on whether the farmer wants the barn along the width (W) or the length (L). If he wants the barn along the width, then the perimeter of the remaining 3 sides equals L+L+W. Subing L=3W gives 3W+3W+W = 240 feet. 7W=240 so W=34.29 feet. And L=3W=102.86 feet. If he wants the barn along the length, then the perimeter of the remaining 3 sides equals W+W+L. Subing L=3W gives W+W+3W = 240 feet. 5W=240 so W=48 feet. And L=3W=144 feet. http://answers.yahoo.com/question/index?qid=20110524125520AAhjrfo

OpenStudy (anonymous):

ok @selenamalter

OpenStudy (ash2326):

We have the condition \[L=3\times W\] Length is three times of width And also the perimeter of the rectangle equals 240 \[2(L+W)=240\] Can you find Length and width from this @selenamalter ?

OpenStudy (anonymous):

144 feet by 48 feet B. 192 feet by 48 feet C. 120 feet by 40 feet D. 108 feet by 58 feet Um....sort of...so it's A

OpenStudy (anonymous):

option a 144 feet by 48 feet

OpenStudy (anonymous):

Right okay